OpenStudy (anonymous):

integrate (x^3) cos 5x dx from x=0 to pi/2

OpenStudy (anonymous):

= 6/625-(3 pi)/125+pi^3/40 = 0.709359

OpenStudy (anonymous):

parts i bet

OpenStudy (anonymous):

indeed. Parts

OpenStudy (anonymous):

Where x^3 = u and cos(5x) = dv

OpenStudy (anonymous):

\[\int\limits _0^{\frac{\pi }{2}}x^3\text{ }\text{Cos}[5 x]dx=\frac{6}{625}-\frac{3 \pi }{125}+\frac{\pi ^3}{40}=0.709359 \]Diogo, are you using Mathemtica?